OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. -334 -274 -284 -346

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@AccessDenied

OpenStudy (raden):

a22 = -226 a1 + (22-1)d = -226 solve for d it is a start

OpenStudy (raden):

so, u have 26 + 21d = -226 move 26 to right right side : 21d = -226 - 26 21d = - 252 d = -252/21 d = 12

OpenStudy (raden):

oppss d= -12 i meant

OpenStudy (raden):

now, to get the value of 31st term : a31 = a1 + (31-1)d a31 = 26 + 30(-12) a31 = 26-360 = ... ?

OpenStudy (anonymous):

-334

OpenStudy (raden):

that's right

OpenStudy (anonymous):

awesome can you assist with one more?

OpenStudy (raden):

post it in the new question, please :)